# Self-Supervised Learning for Robotic Leaf Manipulation: A Hybrid Geometric-Neural Approach

### v1

@@ Abstract
Introduction
Agricultural robotics has emerged as a critical technology for addressing labor shortages and improving efficiency in modern farming operations (Bechar and Vigneault 2016). Among the various tasks in greenhouse cultivation, leaf sampling for disease detection and health monitoring remains a significant bottleneck, requiring skilled workers to manually identify, select, and extract tissue samples from thousands of plants (Shamshiri et al. 2018). This labor-intensive process not only increases operational costs but also limits the frequency and scale of plant health monitoring, potentially allowing diseases to spread undetected (Bac et al. 2014).
Automating leaf manipulation presents unique challenges compared to traditional robotic grasping tasks. Unlike rigid industrial objects, plant leaves are deformable, vary significantly in size and orientation, and are often partially occluded in dense canopies (Lehnert et al. 2017). While recent advances in deep learning have revolutionized robotic grasping for industrial applications (Mahler et al. 2017; Morrison, Corke, and Leitner 2018), these approaches typically require large datasets of labeled grasp points—a resource that is prohibitively expensive to create for agricultural settings where plant morphology varies continuously throughout growth cycles.
Existing approaches to agricultural manipulation fall into two categories: purely geometric methods that rely on hand-crafted features (Hemming et al. 2014a; Silwal et al. 2017), and end-to-end deep learning systems trained on synthetic or limited real-world data (Arad et al. 2020; Yu et al. 2019). Geometric approaches, while interpretable and robust to domain shifts, struggle with the natural variability of plant structures. Conversely, deep learning methods excel at handling complex visual patterns but suffer from poor generalization when deployed on new crop varieties or growth stages not represented in their training data.
We present a novel hybrid approach that leverages the complementary strengths of geometric reasoning and neural networks through self-supervised learning. Our key insight is that traditional computer vision algorithms, despite their limitations, encode valuable domain expertise that can serve as a teacher for training neural networks without manual annotation. This approach enables continuous learning from operational data while maintaining the interpretability and reliability required for agricultural automation.
Our system operates on a 6-DOF gantry robot equipped with stereo vision and a custom end-effector for leaf manipulation. The perception pipeline combines YOLOv8 instance segmentation (Jocher, Chaurasia, and Qiu 2023) with RAFT-Stereo depth estimation (Lipson, Teed, and Deng 2021) to generate 3D representations of plant canopies. For grasp point selection, we implement a dual-path architecture: a geometric pipeline using Pareto optimization across multiple hand-crafted features (flatness, accessibility, edge distance), and a convolutional neural network with spatial attention that learns from the geometric system’s decisions.
@@ Introduction
Related Work
Vision-Based Leaf Manipulation
@@
Traditional approaches to robotic leaf manipulation in agricultural settings relied on geometric reasoning and classical computer vision. Hemming et al. developed methods for cucumber leaf detection in greenhouses using color and texture features (Hemming et al. 2014b), while Bac et al. presented obstacle-aware motion planning for tomato canopies (Bac et al. 2017). Several studies focused on deformable leaf modeling, including Cerutti et al.’s parametric active polygon models (Cerutti et al. 2013) and Xia et al.’s active shape models for overlapping leaves (Xia et al. 2018). The integration of 3D information improved robustness, as demonstrated by Guo and Xu’s multiview stereo reconstruction for lettuce segmentation (Guo and Xu 2017). While effective in controlled conditions, these methods often required extensive tuning and struggled with natural plant variability.
Deep Learning for Agricultural Grasping
@@
Deep learning has shown promise in agricultural manipulation, though with unique challenges compared to industrial applications. Barth et al. developed CNN-based systems for broccoli harvesting that handle significant occlusion (Barth et al. 2019), while Arad et al. demonstrated sweet pepper harvesting combining YOLO detection with stereo depth (Arad et al. 2020). For leaf-specific tasks, Ahlin et al. pioneered CNN-based leaf identification with visual servoing for autonomous sampling, achieving 85% success rates in greenhouses (Ahlin et al. 2016). However, these approaches typically require extensive training data—a significant limitation given the continuous variation in plant morphology. To address this, researchers have explored simulation, with approaches like Dex-Net generating synthetic grasp scenarios (Mahler et al. 2017), inspiring agricultural adaptations for data generation.
Self-Supervised Learning in Agricultural Robotics
@@
Self-supervised learning has emerged as a promising paradigm for agricultural robotics, particularly where manual annotation is expensive. Zhang et al. demonstrated self-supervised learning for tomato harvesting, using classical vision systems to provide training labels (Zhang and Yang 2021). Similar bootstrapping approaches include Kootstra et al.’s work on sweet pepper detection, where geometric algorithms generated training data for neural networks (Kootstra et al. 2021). This knowledge transfer from classical to learning-based systems has proven particularly valuable in controlled environment agriculture, where hybrid approaches consistently outperform purely learned policies (Shamshiri et al. 2018).
3D Perception and Hybrid Systems
@@
Accurate depth sensing is crucial for manipulation in dense plant canopies. While traditional stereo algorithms struggle with plant textures, recent advances like RAFT-Stereo have dramatically improved accuracy for agricultural applications (Lipson, Teed, and Deng 2021). Lipson et al.’s recurrent architecture achieves state-of-the-art performance on challenging plant datasets, enabling precise leaf pose estimation (Sa et al. 2017). Recent research increasingly combines classical and learning approaches, as demonstrated by Lehnert et al.’s hybrid system for pepper harvesting (Lehnert et al. 2016). These hybrid architectures leverage geometric interpretability with neural adaptability, making them ideal for complex agricultural tasks where safety and reliability are paramount (Duckett et al. 2018).
Method
@@
Instance Segmentation
We utilize YOLOv8 (Jocher, Chaurasia, and Qiu 2023) for real-time instance segmentation of individual leaves. Unlike standard implementations, we fine-tuned YOLOv8 on a custom dataset of 900+ images containing soybean and tomato plants in greenhouse conditions. This domain-specific training enables robust leaf detection even in challenging scenarios with significant overlap and occlusion, achieving 90%+ confidence scores in operational conditions.
@@
Stereo Depth Estimation
For 3D reconstruction, we employ RAFT-Stereo (Lipson, Teed, and Deng 2021), which generates dense disparity maps through iterative refinement using recurrent all-pairs field transforms. This approach handles the thin structures and low-texture regions characteristic of plant foliage more reliably than traditional stereo matching algorithms (Scharstein and Szeliski 2002).
@@
Our calibrated stereo pair captures synchronized images at 1440$\times$1080 resolution. As illustrated in Figure 3, RAFT-Stereo processes these to produce sub-pixel accurate disparity maps in approximately 60ms, achieving 29% lower 1-pixel error than previous methods on standard benchmarks (Geiger, Lenz, and Urtasun 2012). The disparity values are converted to metric depth using the camera calibration parameters, enabling accurate 3D reconstruction.
3D Reconstruction
@@
This fusion process outputs a structured representation of each leaf containing both 2D mask information and 3D geometric properties, providing the necessary data for subsequent grasp point selection algorithms. The geometric processing includes signed distance field (SDF) generation, which will be detailed in Section Geometric Feature Scoring Pipeline.
Geometric Feature Scoring Pipeline
@@
Optimal Leaf Selection
@@
Geometric Grasp Point Scoring
@@
Stem Proximity Penalty
@@
Neural Refinement Module (GraspPointCNN)
@@
Self-Supervised Learning Framework
@@
Automatic Training Data Generation
@@
Continuous Learning Pipeline
@@
Hybrid Decision Integration
@@
where $C_{pred}$ is the confidence score described in Section Confidence Estimation. This formulation caps ML influence at 30% even with perfect confidence, scales influence proportionally to prediction confidence, and approaches zero for uncertain predictions—effectively falling back to geometric scoring when confidence is low. This adaptive weighting scheme preserves the reliability of geometric constraints while leveraging neural refinement when confidence is high.
@@
Dataset and Setup
Hardware Configuration
@@
Dataset Collection
The dataset includes tomato (60%) and soybean (40%) plants at various growth stages grown under controlled greenhouse conditions. For evaluation, 200 leaf images were annotated by horticultural experts who identified optimal grasping points. The self-supervised training dataset (875 samples) described in Section Self-Supervised Learning Framework was derived from this collection, while testing used 150 separate stereo image pairs with novel plant arrangements.
Evaluation Metrics
@@
Component Contribution Analysis
@@
Neural Refinement Analysis
@@ Discussion
Comparative Analysis
To evaluate our hybrid approach against existing methods, we conducted comprehensive experiments using the metrics defined in Section Evaluation Metrics.
Baseline Comparison
@@
Comparison to Literature
Our 88.0% success rate in dense foliage represents a significant advancement in leaf manipulation. Ahlin et al. (Ahlin et al. 2016) demonstrated leaf picking using visual servoing but without reporting quantitative success rates. Their monocular approach required careful camera alignment, while our stereo-based system resolves depth ambiguities across varying viewpoints.
For context, robotic fruit harvesting systems typically achieve 70-90% success in less cluttered environments (Silwal et al. 2017; Arad et al. 2020; Bac et al. 2017). Bac et al. (Bac et al. 2017) reported 83% success for sweet pepper harvesting, while Silwal et al. (Silwal et al. 2017) achieved 84% for apples under ideal conditions. Our 88% success in highly cluttered leaf scenarios demonstrates the effectiveness of our approach given the additional challenges of occlusion and thin structures.
Sa et al. (Sa et al. 2017) combined color and 3D information for sweet pepper peduncle detection, achieving 90% detection accuracy but not reporting manipulation success. Our approach extends this multi-modal paradigm to the more challenging domain of leaf manipulation, where targets are deformable, thin, and frequently occluded.
Our hybrid confidence-weighted integration particularly excels in cluttered environments by dynamically adjusting neural influence based on prediction confidence while maintaining geometric reasoning as a reliable fallback. This adaptive integration advances beyond existing agricultural systems that typically rely on either pure geometric reasoning (Bac et al. 2014) or standalone neural approaches (Yu et al. 2019; Ahlin et al. 2016).
Real-World Validation
@@
Qualitative Performance Analysis
@@
The practical validation confirmed that our confidence-weighted approach effectively combines the reliability of geometric constraints with the adaptability of neural refinement, resulting in a robust system capable of autonomous operation in dynamic agricultural environments.
Discussion
@@ Acknowledgments
The development of T-Rex was supported by the USDA-NIFA Cyber-Physical Systems program (Award #2021-67021-34037) to AS and GK at CMU, and SL, BV and CL at Virginia Tech. The metagenomic analysis and machine learning components were supported by USDA FACT-CIN (Award #2021-67021-34343) to SL and BV labs at Virginia Tech. Additional support was provided by the NSF AI Institute for Resilient Agriculture (AIIRA, Award #2021-67021-35329). The authors thank the field robotics team at CMU for their technical support, and collaborators at Virginia Tech for guidance on plant pathogen assays. We would also like to thank Dexter Friis-Hecht, Kalinda Wagner, and Carolin Kiewel for their contributions to the computer vision pipeline, end-effector design, and circuit implementation.

### v3

@@ Abstract
1 Introduction
Agricultural robotics has emerged as a key technology for addressing labor shortages and improving efficiency in modern farming [8, 43]. Among greenhouse cultivation tasks, leaf sampling for disease detection remains a significant bottleneck, requiring skilled workers to manually identify, select, and extract tissue samples from thousands of plants [42, 4]. This labor-intensive process increases operational costs and limits the frequency of plant health monitoring, potentially allowing diseases to spread undetected [5, 33].
Automating leaf manipulation presents unique challenges compared to traditional robotic grasping tasks. Unlike rigid industrial objects, plant leaves are deformable, vary significantly in size and orientation, and are often partially occluded in dense canopies [30, 38]. While recent advances in deep learning have revolutionized robotic grasping for industrial applications [34, 37, 53], these approaches typically require large datasets of labeled grasp points—a resource that is prohibitively expensive to create for agricultural settings where plant morphology varies continuously throughout growth cycles [26].
Existing approaches to agricultural manipulation fall into two categories: purely geometric methods that rely on hand-crafted features [18, 44, 20], and end-to-end deep learning systems trained on synthetic or limited real-world data [3, 52, 49]. Geometric approaches, while interpretable and robust to domain shifts, struggle with the natural variability of plant structures. Conversely, deep learning methods excel at handling complex visual patterns but suffer from poor generalization when deployed on new crop varieties or growth stages not represented in their training data [36].
We present a novel hybrid approach that leverages the complementary strengths of geometric reasoning and neural networks through self-supervised learning. Our key insight is that traditional computer vision algorithms, despite their limitations, encode valuable domain expertise that can serve as a teacher for training neural networks without manual annotation [21]. This approach enables continuous learning from operational data while maintaining the interpretability and reliability required for agricultural automation.
Our system operates on a 6-DOF gantry robot equipped with stereo vision and a custom end-effector for leaf manipulation. The perception pipeline combines YOLOv8 instance segmentation [23] with RAFT-Stereo depth estimation [31] to generate 3D representations of plant canopies. For grasp point selection, we implement a dual-path architecture: a geometric pipeline using Pareto optimization across multiple hand-crafted features (flatness, accessibility, edge distance), and a convolutional neural network with spatial attention that learns from the geometric system’s decisions [25].
@@ 1 Introduction
2 Related Work
2.1 Vision-Based Leaf Manipulation
@@
Traditional approaches to robotic leaf manipulation in agricultural settings relied on geometric reasoning and classical computer vision. Hemming et al. developed methods for cucumber leaf detection in greenhouses using color and texture features [19], while Bac et al. presented obstacle-aware motion planning for tomato canopies [6]. Several studies focused on deformable leaf modeling, including Cerutti et al.’s parametric active polygon models [9], Xia et al.’s active shape models for overlapping leaves [51], and Jin et al.’s probabilistic graphical models for leaf structure analysis [22]. The integration of 3D information improved robustness, as demonstrated by Guo and Xu’s multiview stereo reconstruction for lettuce segmentation [16] and Sodhi et al.’s plant growth monitoring system using structure-from-motion [45]. While effective in controlled conditions, these methods often required extensive tuning and struggled with natural plant variability, particularly under varying illumination conditions [11].
2.2 Deep Learning for Agricultural Grasping
@@
Deep learning has shown promise in agricultural manipulation, though with unique challenges compared to industrial applications. Barth et al. developed CNN-based systems for broccoli harvesting that handle significant occlusion [7], while Arad et al. demonstrated sweet pepper harvesting combining YOLO detection with stereo depth [3]. For leaf-specific tasks, Ahlin et al. pioneered CNN-based leaf identification with visual servoing for autonomous sampling, achieving 85% success rates in greenhouses [2]. However, these approaches typically require extensive training data—a significant limitation given the continuous variation in plant morphology [47]. To address this, researchers have explored simulation, with approaches like Dex-Net generating synthetic grasp scenarios [34], inspiring agricultural adaptations for data generation [28]. Recent advances in generative data augmentation have shown promise for bridging the synthetic-real domain gap [48], particularly for multi-crop environments with challenging lighting conditions.
2.3 Self-Supervised Learning in Agricultural Robotics
@@
Self-supervised learning has emerged as a promising paradigm for agricultural robotics, particularly where manual annotation is expensive. Zhang et al. demonstrated self-supervised learning for tomato harvesting, using classical vision systems to provide training labels [54]. Similar bootstrapping approaches include Kootstra et al.’s work on sweet pepper detection, where geometric algorithms generated training data for neural networks [27] and Tao et al.’s approach using temporal consistency for plant growth tracking [46]. This knowledge transfer from classical to learning-based systems has proven particularly valuable in controlled environment agriculture, where hybrid approaches consistently outperform purely learned policies [42, 14]. Recent work has also explored contrastive learning frameworks for agricultural visual representations [50], enabling more sample-efficient adaptation to new crops and growth conditions.
2.4 3D Perception and Hybrid Systems
@@
Accurate depth sensing is crucial for manipulation in dense plant canopies. While traditional stereo algorithms struggle with plant textures, recent advances like RAFT-Stereo have dramatically improved accuracy for agricultural applications [31]. Lipson et al.’s recurrent architecture achieves state-of-the-art performance on challenging plant datasets, enabling precise leaf pose estimation [40]. Alternative sensing modalities such as time-of-flight cameras [35] and structured light systems [55] have also shown promise for plant phenotyping applications with complex geometries. Recent research increasingly combines classical and learning approaches, as demonstrated by Lehnert et al.’s hybrid system for pepper harvesting [29] and Adamides et al.’s framework for integrating human expertise with machine learning [1]. These hybrid architectures leverage geometric interpretability with neural adaptability, making them ideal for complex agricultural tasks where safety and reliability are paramount [12]. Confidence-aware systems that adaptively balance multiple decision sources have proven especially effective in scenarios with high uncertainty, demonstrating robustness across varying environmental conditions [13].
3 Method
@@ 3.2 Vision Pipeline
3.2.1 Instance Segmentation
We utilize YOLOv8 [23] for real-time instance segmentation of individual leaves. Unlike standard implementations, we fine-tuned YOLOv8 on a custom dataset of 900+ images containing soybean and tomato plants in greenhouse conditions. This domain-specific training enables robust leaf detection even in challenging scenarios with significant overlap and occlusion, achieving 90%+ confidence scores in operational conditions.
@@ 3.2.1 Instance Segmentation
3.2.2 Stereo Depth Estimation
For 3D reconstruction, we employ RAFT-Stereo [31], which generates dense disparity maps through iterative refinement using recurrent all-pairs field transforms. This approach handles the thin structures and low-texture regions characteristic of plant foliage more reliably than traditional stereo matching algorithms [41].
@@ 3.2.2 Stereo Depth Estimation
Our calibrated stereo pair captures synchronized images at 1440$\times$1080 resolution. As illustrated in Figure 3, RAFT-Stereo processes these to produce sub-pixel accurate disparity maps in approximately 60ms, achieving 29% lower 1-pixel error than previous methods on standard benchmarks [15]. The disparity values are converted to metric depth using the camera calibration parameters, enabling accurate 3D reconstruction.
3.2.3 3D Reconstruction
@@ 3.2.4 Data Fusion
This fusion process outputs a structured representation of each leaf containing both 2D mask information and 3D geometric properties, providing the necessary data for subsequent grasp point selection algorithms. The geometric processing includes signed distance field (SDF) generation, which will be detailed in Section 3.3.
3.3 Geometric Feature Scoring Pipeline
@@
3.3.1 Optimal Leaf Selection
@@
3.3.2 Geometric Grasp Point Scoring
@@
3.3.3 Stem Proximity Penalty
@@
3.4 Neural Refinement Module (GraspPointCNN)
@@ 3.4.3 Confidence Estimation
3.5 Self-Supervised Learning Framework
@@
3.5.1 Automatic Training Data Generation
@@ 3.5.2 Training Methodology
3.5.3 Continuous Learning Pipeline
@@
3.6 Hybrid Decision Integration
@@
where $C_{pred}$ is the confidence score described in Section 3.4.3. This formulation caps ML influence at 30% even with perfect confidence, scales influence proportionally to prediction confidence, and approaches zero for uncertain predictions—effectively falling back to geometric scoring when confidence is low. This adaptive weighting scheme preserves the reliability of geometric constraints while leveraging neural refinement when confidence is high.
@@ 4 Experiments and Results
4.1 Dataset and Setup
4.1.1 Hardware Configuration
@@
4.1.2 Dataset Collection
The dataset includes tomato (60%) and soybean (40%) plants at various growth stages grown under controlled greenhouse conditions. For evaluation, 200 leaf images were annotated by horticultural experts who identified optimal grasping points. The self-supervised training dataset (875 samples) described in Section 3.5 was derived from this collection, while testing used 150 separate stereo image pairs with novel plant arrangements.
4.1.3 Evaluation Metrics
@@ 4.2 Ablation Studies
4.2.1 Component Contribution Analysis
@@
4.2.2 Neural Refinement Analysis
@@ 4.2.3 Discussion
4.3 Comparative Analysis
To evaluate our hybrid approach against existing methods, we conducted comprehensive experiments using the metrics defined in Section 4.1.
4.3.1 Baseline Comparison
@@
4.3.2 Comparison to Literature
Our 88.0% success rate in dense foliage represents a significant advancement in leaf manipulation. Ahlin et al. [2] demonstrated leaf picking using visual servoing but without reporting quantitative success rates. Their monocular approach required careful camera alignment, while our stereo-based system resolves depth ambiguities across varying viewpoints, similar to approaches that explicitly model uncertainty in depth perception [10].
For context, robotic fruit harvesting systems typically achieve 70-90% success in less cluttered environments [44, 3, 6]. Bac et al. [6] reported 83% success for sweet pepper harvesting, while Silwal et al. [44] achieved 84% for apples under ideal conditions. Kang et al. [24] emphasized the importance of standardized metrics for agricultural manipulation, noting that success rates for thin, deformable targets typically lag 10-15% behind rigid object grasping. Our 88% success in highly cluttered leaf scenarios demonstrates the effectiveness of our approach given the additional challenges of occlusion and thin structures, exceeding the performance bounds established in previous comparative studies [39].
Sa et al. [40] combined color and 3D information for sweet pepper peduncle detection, achieving 90% detection accuracy but not reporting manipulation success. Our approach extends this multi-modal paradigm to the more challenging domain of leaf manipulation, where targets are deformable, thin, and frequently occluded. Recent work by Liu et al. [32] on deformable leaf modeling achieved 78% grasping success but required significantly longer planning times (350-450ms) compared to our 158.7ms.
Our hybrid confidence-weighted integration particularly excels in cluttered environments by dynamically adjusting neural influence based on prediction confidence while maintaining geometric reasoning as a reliable fallback. This adaptive integration advances beyond existing agricultural systems that typically rely on either pure geometric reasoning [5] or standalone neural approaches [52, 2]. Similar confidence-aware fusion strategies have shown promising results in medical robotics [17], though with significantly higher computational requirements that limit real-time performance in field conditions.
4.4 Real-World Validation
@@ 4.4.1 Operational Deployment
4.4.2 Qualitative Performance Analysis
@@
The practical validation confirmed that our confidence-weighted approach effectively combines the reliability of geometric constraints with the adaptability of neural refinement, resulting in a robust system capable of autonomous operation in dynamic agricultural environments. “‘
5 Discussion
@@ Acknowledgments
The development of T-Rex was supported by the USDA-NIFA Cyber-Physical Systems program (Award #2021-67021-34037) and USDA FACT-CIN (Award #2021-67021-34343). Additional support was provided by the NSF AI Institute for Resilient Agriculture (AIIRA, Award #2021-67021-35329). The authors thank the field robotics team at CMU for their technical support and collaborators at Virginia Tech for guidance on plant pathogen assays. We would also like to thank Dexter Friis-Hecht, Kalinda Wagner, and Carolin Kiewel for their contributions to the computer vision pipeline, end-effector design, and circuit implementation.